OpenStudy (anonymous):

5(2x - 6) = 2x + 2 solve for X. I tried doing it like this: 5(2x - 6) = 2x + 2 10x - 6 = 2x 15x - 8 = x but then I got stuck, because I decided the last thing I did didn't get rid of anything. Help?

OpenStudy (anonymous):

just realized I typed a number wrong. just edited it

OpenStudy (anonymous):

The left side of your equation can be distributed. 5(2x-6) = 10x-30 10x-30=2x+2 Then, you're going to want to move all terms containing "x" to the left, and all the other terms to the right. 10x-30-2x=2 Combining like terms, 10x - 2x = 8x. -30+8x = 2 Then, you're going to add 30 to both sides of the equation. 8x=2 + 30 8x = 32 Divide each side by 8. x = 4 And there's your answer. Does that make sense?

OpenStudy (anonymous):

i think that makes sese yeah

OpenStudy (anonymous):

X=4 is the answer

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you well come my friend